Shannon had three times as many quarters as nickels. She had $2.40 in all. How many quarters did she have?
Answer by stanbon(75887)

Shannon had three times as many quarters as nickels. She had $2.40 in all. How many quarters did she have?
---
Equations:
q = 3n
5n + 25q = 240
----
Substitute for q and solve for n:
5n + 25(3n) = 240
5n + 75n = 240
80n = 240
n = 3 (number of nickels)
----
Since q = 3n, q = 9 (number of quarters)
=========
Cheers,
Stan H.